A machine that manufactures automobile parts is estimated to produce defective parts 8% of the time. If this estimate is correct, and 10 parts produced by this machine are randomly selected, what is the probability that more than 1 turn out to be defective? Round your answer to four decimal places.

Answer by stanbon(75887)

A machine that manufactures automobile parts is estimated to produce defective parts 8% of the time. If this estimate is correct, and 10 parts produced by this machine are randomly selected, what is the probability that more than 1 turn out to be defective? Round your answer to four decimal places.
----------------------------------------
It's a binomial problem with n=10, p = 0.08 and x>=2
----------------------------------------------------
P(x>=2) = 1 -[P(x=0) + P(x=1)]
P(x>=2) = 1 -[(0.98)^10 + 10*(0.98)^9(0.02)] = 0.016776..
============================================================
Cheers,
Stan H.